OpenStudy (anonymous):

can i solve \[\lim_{x \rightarrow \infty} (2-x)/\sqrt(9x^2 +1)\] without using l'hopitals rule?

OpenStudy (anonymous):

sure you can do it with your eyeballs

OpenStudy (anonymous):

you mean algebra?

OpenStudy (anonymous):

numerator is like -x, denominator is like 3x and \[\frac{-x}{3x}=-\frac{1}{3}\]

OpenStudy (anonymous):

no i meant eyeballs. just by looking

OpenStudy (ash2326):

yeah take x^2 out of the square root \[(2-x) / (x \sqrt{9+1/x^2})\] take x out from numerator \[x(2/x-1) / (x \sqrt{9+1/x^2})\] \[(2/x-1) / ( \sqrt{9+1/x^2})\] 1/x and 1/x^2 ---->0 as x---> \(\infty\) so we've \[ -1/ \sqrt 9\] -1/3

OpenStudy (anonymous):

that is ok, but seems like a lot of work

OpenStudy (anonymous):

thanks ash for explanation, satellite your wit is endless

OpenStudy (anonymous):

but not helpful

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

actually i was quite serious. suppose you wanted \[\lim_{x\rightarrow \infty}\frac{2x^2+3x+1}{5x^2+x-2}\] wouldn't you say \[\frac{2}{5}\] just like that?